three capacitors c1=20mf  c2=40mf   c3=50mf are connected in series .if no capacitor can sustain more than 50v ,then the maximum potential difference between the two terminals X and Y is?

Grade:12

2 Answers

Arun Kumar IIT Delhi
9 years ago
In series charges are same.

We get Q from here then we will get V by RHS.

Thanks & Regards
Arun Kumar
IIT Delhi
Askiitians Faculty
Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem below.
 
Let, the charge on each capacitor be q.
Now, Potential drop across a capacitor, V = Q/C
Hence, the potential drop across the given capacitors will be-
For C1, V1 = q/20
For C2, V2 = q/40
For C3, V3 = q/50
 
Now, V1 is the highest among all the individual voltages.
Its max value can reach upto 50 volts
Hence, q/20 = 50
or, q = 1000 mC
 
Hence, V1 = 50 V
V2 = 1000/40 = 25 V
V3 = 1000/50 = 20 V
 
Hence potential difference between x and y terminal, V  = V1 + V2 + V3
                                                                             = 50 + 25 + 20
                                                                             = 95 V
 
Thanks and regards,
Kushagra